- Two small silver spheres each with a mass of 100gm are separated by 1 m. Calculate the number of electrons that must be transferred from one to the other to produce an attractive force of 104 N between the spheres.


1
Expert's answer
2021-12-27T08:01:24-0500

Apply Coulomb's law:


"F=\frac1{4\pi\epsilon_0}\frac{(ne)^2}{r^2},\\\space\\
n=\frac {2r}e\sqrt{\pi\epsilon_0F}=6.58·10^{15}"

assuming the force is 104 N, not 104 N.